OpenStudy (anonymous):

What is the slope of the line 5x + 2y = 10?

OpenStudy (anonymous):

set it to y-int form

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

where m is the slope

OpenStudy (anonymous):

that means isolating y in the given equation

OpenStudy (anonymous):

Still confused? ask more questions.

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

thanks for giving the answer away

OpenStudy (anonymous):

\[2y = -5x + 10\] \[y = -5x/2 + 10/2\] \[y = -5x/2 + 5\] Comparing with y = mx + c, We get m = -5/2..